# Implicit Geometric Descriptor-Enabled ANN Framework for a Unified Structure-Property Relationship in Architected Nanofibrous Materials

### v1

@@ 1 Introduction
To address this challenge, we introduce an artificial neural network (ANN) framework that correlates two multi-component shape invariant (MCSI) implicit geometric descriptors (Fig. 1(a)(gray center) and Fig. 1(d)) [22, 23] of the architecture to the experimentally measured specific mechanical properties, such as specific energy absorption ($W^{*}$), specific peak stress ($\sigma^{*}$), and specific average modulus ($E^{*}$) to establish a unified structure-property relationship. These implicit geometric descriptors capture the size of the features within the architecture and the compactness of the architecture, which give rise to the nanoscale mechanisms that govern the bulk mechanical properties (Fig. 1 (b-c)). Our implicit geometric descriptor-based approach enables both the prediction of mechanical properties of new architectures and the design of architectures for diverse mechanical property combinations through independent control of these descriptors. The findings of this study accentuates that the nanoscale mechanisms—such as size effects and inter-nanotube interactions—predominantly govern the mechanical properties of architected VACNT foams, rather than the specific spatial architecture. Unlike the complex deep learning models with numerous parameters, the reduced parameter space enabled by the MCSI implicit descriptors allows faster model training with much fewer experimental measurements—enabling efficient and accelerated design of architected materials where such multi-scale mechanisms interactively give rise to bulk mechanical behavior.
@@ 2.1 Synthesis of 2D architected VACNT foams
We synthesized the architected VACNT foams using floating catalyst thermal chemical vapor deposition (tCVD) [24] process on patterned p-type silicon wafer with (100) crystal orientation. First, we spin coated the substrate with S1813 photoresist at 3000 rpm for 30 seconds to a thickness of 10 $\mu m$. Then, we partially diced the 100 $mm$ diameter wafer to 30% of its thickness (500 $\mu m$), creating a grid of 5 $mm$$\times$ 5 $mm$ squares (A CAD file created in AutoCAD defined a grid of 5 $mm$$\times$ 5 $mm$ squares on 4-inch substrate with each square containing the desired micropattern). We transfered the desired micropattern on to the photoresist-coated substrate using a Heidelberg DWL 66+ laser writer, which utilizes a 405 $nm$ UV laser diode for the process. Initially, the CAD file was converted into DXF format using “x-convert” software and imported into the DWL 66 interface. We carefully mounted the coated substrate on the system’s holder while ensuring accurate alignment and leveling. A 10 $mm$ write head, capable of achieving a minimum feature size of  1 $\mu m$, was used for the exposure. We calibrated the parameters such as laser power (60 $mW$), intensity (100%), and filter (100%) based on the properties of S1813 photoresist and silicon substrate. After exposure, the we removed the unexposed photoresist by developing the wafer in an MF321 developer bath for 30 seconds. Next, we evaporated a chromium layer of 20 $nm$ at a rate of 0.5 Å/$s$ onto the wafer using a metal evaporator. We then removed the UV-exposed photoresist in an acetone bath, leaving a chromium film with the inverse pattern of the desired VACNT architecture on the substrate that masked predefined areas preventing CNT growth.
@@
All samples had a constant areal dimension of $5$ $mm$$\times$ $5$ $mm$ with heights ($h$) ranging from 1 to 3 $mm$ (Tables S1-S3). We designed 81 different architectures that includes 60 cylindrical architectures [20], 18 concentric cylindrical architectures [21], and 3 fractal architectures [11]. We designed the 60 cylindrical architectures using a full factorial design of experiments (DOE) approach, varying the explicit geometric design parameters, $D_{in}$, $t$ and $g$. To create concentric cylindrical architectures, explicit design parameters, $D_{in}$, $t$, $g_{cc}$ and $n$ were varied, keeping the $D_{in}/t$ ratio at 5, while for fractal arrays, $m$ was varied from 1 to 3. Table 1 summarizes the different levels of explicit design parameters used to design the cylindrical, concentric cylindrical, and the fractal arrays.
To test and validate our ANN models on designs that go beyond these regular cylinder-based architectures described by their respective explicit geometric design parameters, we designed additional cylinder-based architectures (Fig. 5(a-d)) and sinusoid-based architectures (Fig. 5(e-i))) that have different explicit geometric design parameters. We measured the mass of individual samples using analytical micro-balance and calculated the volume from dimensions measured using precision calipers ($5mm\times 5mm\times h$) (Tables S1-S3 and Table S4). We then determined the bulk density of the architected samples ($\rho$) by dividing their mass by the volume.
@@
Figure 2: Variation of implicit geometric descriptors—$\mathcal{I}_{mc-c}$ (cyan) and $\mathcal{D}$ (violet)—across different architectures as a function of their explicit design parameters (a) $D_{in}$ in cylindrical architecture, (b) $g$ in cylindrical architecture, (c) $t$ in cylindrical architecture, (d) $g_{cc}$ in concentric cylindrical architecture, (e) $n$ in concentric cylindrical architecture and (f) $m$ in fractal arrays. (Top panels illustrate the explicit design parameters)
@@ 2.2 Mechanical testing of architected VACNT foams and calculation of their specific mechanical properties
We measured the mechanical properties of all the architected VACNT foams used in this work using quasistatic uniaxial compression experiments [20, 21, 11]. We performed these experiments using an Instron E3000 Electropulse universal testing system. The system was equipped with a 5 $kN$ load cell and was controlled through the Instron software, “wavematrix”. We controlled the maximum compressive strain and the strain rate by prescribing the displacement and displacement rate using the wavematrix. The specific mechanical properties of interest—$W^{*}$, $\sigma^{*}$ and $E^{*}$—were calculated from the quasistatic compressive stress-strain curves and the bulk densities of the architected samples using Equations 1, 2, and 3 respectively:
@@ 3.2 Development of the ANN framework to establish unified structure-property correlations
Leven-Marquardt (LM) [36] and Bayesian regularization (BR) [37] are the commonly used optimization techniques to train ANNs with relatively small datasets. While LM is often used for faster convergence in nonlinear least-squares problems, it lacks emphasis on regularization, which is important for avoiding overfitting, especially with limited data. To enhance the model’s generalization and robustness to unseen data, we opted for the BR backpropagation algorithm, which reduces model complexity by regularizing the cost function [38, 39, 40]. Unlike other early stopping-based methods, this method does not require a validation set to monitor performance [40], instead it incorporates prior knowledge about weight distribution and penalizes complex models for unreasonable assignment of weights, leading to better generalization. We used the mean square error (MSE) (Equation 11) as the cost function:
@@
Here, $Y_{i}$ and $\hat{Y}_{i}$ denote the predicted and target output variables, whereas $N$ is the number of observations. There was no need to manually set a threshold for the cost function, as BR automatically adjusts parameters like regularization terms and stopping criteria based on the data and model complexity.
@@
Figure 5: Generalizability of ANN models to cylinder-based and sinusoid-based designs. SEM images of cylinder-based fractal designs created by incorporating concentric rings, with corresponding unit cells shown below: (a) $C_{1}$ $1^{st}$ order fractal array with concentric rings within the inscribed cylinders, (b) $C_{2}$ $1^{st}$ order fractal array with concentric rings within both the inscribed and circumscribed cylinders, (c) $C_{3}$ $2^{nd}$ order fractal array with concentric rings within $1^{st}$ order inscribed cylinders and (d) $C_{4}$ $2^{nd}$ order fractal array with concentric rings within all inscribed and the circumscribed cylinders. SEM images of sinusoid-based designs with corresponding sine function, thickness and diameter of the inscribed cylinder given below: (e) $S_{1}$ One-way sinusoid pattern, (f) $S_{2}$ One-way sinusoid pattern with cylinders applied within gaps, (g) $S_{3}$ Two-way sinusoid pattern, (h) $S_{4}$ Two-way sinusoid pattern with a different sine function and (i) $S_{5}$ Two-way sinusoid pattern with cylinders applied within gaps. Comparison of values predicted by ANN with the experimentally measured values across architectures $C_{1}$ to $S_{5}$ for (j)$W^{*}$, (k) $\sigma^{*}$ and (l) $E^{*}$. (m) Ratio ($\eta$) of ANN-predicted to the experimentally-measured values for $W^{*}$, $\sigma^{*}$ and $E^{*}$ across $C_{1}$ to $S_{5}$
@@
For each mechanical property, the ANN differed only in the optimal number of neurons in the hidden layer. The grid search-based parameter optimization process (Fig. S1) that we employed is illustrated in Supplementary Information . We divided the input dataset of 81 designs into two subsets, 66 designs for training, and 15 designs for testing. Using a holdout validation approach (shuffle-split procedure) [44, 42], we trained the network multiple times with different shuffled splits, then averaged the performance metrics to obtain a more reliable estimate of the model’s overall performance.
The ANN model demonstrated excellent correlations for all three mechanical properties under consideration, indicating its effectiveness in capturing the structure-property relationships. As shown in Fig. 4(b-g), the coefficient of determination ($R^{2}$) greater than 0.8 and mean absolute percentage errors (MAPE) around 10-15 % (Table 2) for each property highlight the model’s impressive performance after training. These results underscore the capability of the implicit geometric descriptor-based ANN approach in successfully establishing a unified structure-property relationship that integrates various types of 2D architectures. We compared the performance of our ANN models with other regression models (see Supplementary Information), including multivariate linear regression (MVLR), MVLR with an added interdependence term given by the product of the descriptors (MVLR-I) (Equation S1) and Gaussian process regression (GPR) (Equation S2). Table 2 provides the summary of the $R^{2}$ and MAPE obtained with each model and demonstrates that the ANN models outperform other regression models in establishing accurate correlations. The relatively strong performance of other regression models also highlight the effectiveness of the implicit geometric descriptor-based approach in capturing the hidden structure-property correlations. Our ANN’s ability to establish a unified structure-property relationship using implicit geometric descriptors that capture the size effect and interaction effect suggests that, in architected VACNT foams, mechanisms like size effect and interaction effect are the primary drivers of bulk mechanical properties, irrespective of the specific 2D architecture—a finding that underscores the utility of the geometric descriptors in ANN rather than mere image regression-based approaches.
@@
Figure 6: Design of architected materials with implicit geometric descriptors. (a) Schematic illustration of the various training conditions used to evaluate the descriptor importance scores—$\Omega_{\mathcal{I}_{mc-c}}$ and $\Omega_{\mathcal{D}}$—and the coefficients of determination—$R^{2}_{\mathcal{I}_{mc-c}}$ and $R^{2}_{\mathcal{D}}$—for the descriptors $\mathcal{I}_{mc-c}$ and $\mathcal{D}$, respectively. (b) Comparison of the descriptor importance scores—$\Omega_{\mathcal{I}_{mc-c}}$ and $\Omega_{\mathcal{D}}$—demonstrating the influence of each descriptor on specific mechanical properties $W^{*}$, $\sigma^{*}$, and $E^{*}$; lower values of $\Omega_{\mathcal{D}}$ in spite of $\Omega_{\mathcal{I}_{mc-c}}$ not changing significantly implies the interdependence between the descriptors in influencing $\sigma^{*}$ and $E^{*}$. (c) Comparison of the coefficients of determination—$R^{2}_{\mathcal{I}_{mc-c}}$ and $R^{2}_{\mathcal{D}}$—showing similar trend as in the case of descriptor importance scores. Design contours for the following specific mechanical properties as functions of implicit geometric descriptors predicted by the ANN model for (d) $W^{*}$, (e) $\sigma^{*}$ and (f) $E^{*}$

### v2

@@ 1 Introduction
To address this challenge, we introduce an artificial neural network (ANN) framework that correlates two multi-component shape invariant (MCSI) implicit geometric descriptors (see section 3.1) (Fig. 1(a)(gray center) and Fig. 1(d)) [22, 23] of the architecture to the experimentally measured specific mechanical properties, such as specific energy absorption ($W^{*}$), specific peak stress ($\sigma^{*}$), and specific average modulus ($E^{*}$) to establish a unified structure-property relationship. These implicit geometric descriptors capture the size of the features within the architecture and the compactness of the architecture, which give rise to the nanoscale mechanisms that govern the bulk mechanical properties (Fig. 1 (b-c)). Our implicit geometric descriptor-based approach enables both the prediction of mechanical properties of new architectures and the design of architectures for diverse mechanical property combinations through independent control of these descriptors. The findings of this study accentuates that the nanoscale mechanisms—such as size effects and inter-nanotube interactions—predominantly govern the mechanical properties of architected VACNT foams, rather than the specific spatial architecture. Unlike the complex deep learning models with numerous parameters, the reduced parameter space enabled by the MCSI implicit descriptors allows faster model training with much fewer experimental measurements—enabling efficient and accelerated design of architected materials where such multi-scale mechanisms interactively give rise to bulk mechanical behavior.
@@ 2.1 Synthesis of 2D architected VACNT foams
We synthesized the architected VACNT foams using floating catalyst thermal chemical vapor deposition (tCVD) [24] process on patterned p-type silicon wafer with (100) crystal orientation. First, we spin coated the substrate with S1813 photoresist at 3000 rpm for 30 seconds to a thickness of 10 $\mu m$. Then, we partially diced the 100 $mm$ diameter wafer to 30% of its thickness (500 $\mu m$), creating a grid of 5 $mm$$\times$ 5 $mm$ squares (A CAD file created in AutoCAD defined a grid of 5 $mm$$\times$ 5 $mm$ squares on 4-inch substrate with each square containing the desired micropattern). We transfered the desired micropattern on to the photoresist-coated substrate using a Heidelberg DWL 66+ laser writer, which utilizes a 405 $nm$ UV laser diode for the process. Initially, the CAD file was converted into DXF format using “x-convert” software and imported into the DWL 66 interface. We carefully mounted the coated substrate on the system’s holder while ensuring accurate alignment and leveling. A 10 $mm$ write head, capable of achieving a minimum feature size of  1 $\mu m$, was used for the exposure. We calibrated the parameters such as laser power (60 $mW$), intensity (100%), and filter (100%) based on the properties of S1813 photoresist and silicon substrate. After exposure, we removed the unexposed photoresist by developing the wafer in an MF321 developer bath for 30 seconds. Next, we evaporated a chromium layer of 20 $nm$ at a rate of 0.5 Å/$s$ onto the wafer using a metal evaporator. We then removed the UV-exposed photoresist in an acetone bath, leaving a chromium film with the inverse pattern of the desired VACNT architecture on the substrate that masked predefined areas preventing CNT growth.
@@
All samples had a constant areal dimension of $5$ $mm$$\times$ $5$ $mm$ with heights ($h$) ranging from 1 to 3 $mm$ (Tables S2-S4). We designed 81 different architectures that includes 60 cylindrical architectures [20], 18 concentric cylindrical architectures [21], and 3 fractal architectures [11]. We designed the 60 cylindrical architectures using a full factorial design of experiments (DOE) approach, varying the explicit geometric design parameters, $D_{in}$, $t$ and $g$. To create concentric cylindrical architectures, explicit design parameters, $D_{in}$, $t$, $g_{cc}$ and $n$ were varied, keeping the $D_{in}/t$ ratio at 5, while for fractal arrays, $m$ was varied from 1 to 3. Table 1 summarizes the different levels of explicit design parameters used to design the cylindrical, concentric cylindrical, and the fractal arrays.
To test and validate our ANN models on designs that go beyond these regular cylinder-based architectures described by their respective explicit geometric design parameters, we designed additional cylinder-based architectures (Fig. 5(a-d)) and sinusoid-based architectures (Fig. 5(e-i))) that have different explicit geometric design parameters. We measured the mass of individual samples using analytical micro-balance and calculated the volume from dimensions measured using precision calipers ($5mm\times 5mm\times h$) (Tables S2-S4 and Table S5). We then determined the bulk density of the architected samples ($\rho$) by dividing their mass by the volume.
@@
Figure 2: Variation of implicit geometric descriptors—$\mathcal{I}_{mc-c}$ (cyan) and $\mathcal{D}$ (violet)—across different architectures as a function of their explicit design parameters (a) $D_{in}$ in cylindrical architecture, (b) $g$ in cylindrical architecture, (c) $t$ in cylindrical architecture, (d) $g_{cc}$ in concentric cylindrical architecture, (e) $n$ in concentric cylindrical architecture and (f) $m$ in fractal arrays. (g-i) Representative specific stress-strain profiles for each architecture type, with shaded regions indicating standard deviation in the measured specific stress: (g) for cylindrical [20], (h) for concentric cylindrical [21] (i) for fractal array [11] (Top panels illustrate the explicit design parameters)
@@ 2.2 Mechanical testing of architected VACNT foams and calculation of their specific mechanical properties
We measured the mechanical properties of all the architected VACNT foams used in this work using uniaxial compression experiments at a quasistatic strain rate of 0.01 $s^{-1}$[20, 21, 11]. We performed these experiments using an Instron E3000 Electropulse universal testing system—designed to accommodate soft materials—employing platens and fixtures carefully matched to both machine and sample compliance, as well as to the specific sample geometry. The system was equipped with a 5 $kN$ load cell and was controlled through the Instron software, “wavematrix”. We controlled the maximum compressive strain and the strain rate by prescribing the displacement and displacement rate using the wavematrix. The specific mechanical properties of interest—$W^{*}$, $\sigma^{*}$ and $E^{*}$—were calculated from the quasistatic compressive stress-strain curves (Fig. 2(g-i) and the bulk densities of the architected samples using Equations 1, 2, and 3 respectively:
@@
Equation 1 for $W^{*}$ represents the area under the loading-unloading hysteresis loop, while Equation 2 for $\sigma^{*}$ represents the stress measured at $\epsilon_{max}$. The $E^{*}$ in Equation 3 is calculated by integrating the mean stress $\tilde{\sigma}$—averaged over the loading and unloading branches—up to $\epsilon_{max}$ and then normalizing by $\epsilon_{max}$. We adopt this integral modulus, rather than the conventional tangent modulus calculated from the initial loading curve, for two key reasons. First, architected VACNT foams, especially those designed for protective applications, typically operate under large deformations where the stress–strain response is nonlinear and continuously evolving. More importantly certain coarse architectures—such as concentric cylinders with large $g_{cc}$—undergo transition from shell buckling to a structural buckling mode, introducing sub‑linear trends (Fig. 5(j-l)) in the stress-strain response at intermediate strains [21]. This response further evolves at higher strains where the nanoscale mechanisms such as size effect and nanoscale interactions become more dominant. Therefore, an average modulus estimated over a large deformation range offers a more representative measure of the foam’s overall response relevant to protective applications. Second, VACNT foams primarily dissipate energy through friction arising from inter-nanotube interactions. This affects not only the loading behavior but also the unloading behavior—that also contributes to the material’s dissipation of the stored elastic energy. Therefore, by calculating the average modulus using the mean stress over both loading and unloading segments, we capture a more accurate representation of the foam’s elastic energy storage [21].
@@ 3 Results and discussion
Explicit geometric parameters are inherently architecture‑dependent: for example, $D_{in}$ is well defined for cylinder-based designs where it explicitly describes the inner diameter of cylinders, however it’s meaning becomes ambiguous in non‑cylindrical designs such as sinusoidal and other arbitrary designs. In addition, each such explicit geometric parameter describes only an isolated geometric feature of the architecture, thereby omitting the collective geometric attributes that primarily govern the general bulk mechanical behavior across various architectures. These limitations hinder the development of a truly universal structure–property relationship for architected VACNT foams. By contrast, implicit geometric descriptors that are both architecture‑agnostic and describe global features of the architecture that give rise to the key phenomena—universally governing the bulk mechanical behavior across various architectures—provide a more robust basis to model the unified structure-property relationship.
@@ 3.2 Development of the ANN framework to establish unified structure-property correlations
Leven-Marquardt (LM) [36] and Bayesian regularization (BR) [37] are the commonly used optimization techniques to train ANNs with relatively small datasets. While LM is often used for faster convergence in nonlinear least-squares problems, it lacks emphasis on regularization, which is important for avoiding overfitting, especially with limited data (Figure S2). To enhance the model’s generalization and robustness to unseen data, we opted for the BR backpropagation algorithm, which reduces model complexity by regularizing the cost function [38, 39, 40]. Unlike other early stopping-based methods, this method does not require a validation set to monitor performance [40], instead it incorporates prior knowledge about weight distribution and penalizes complex models for unreasonable assignment of weights, leading to better generalization (see Supplementary Information and Figure S2). Therefore, BR is also a generally preferred regularization technique— particularly with smaller datasets—over traditional regularization techniques such as L1 and L2 regularization methods, that rely on cross‑validated hyperparameters and dropout methods that often underfit with small datasets [41, 42].
We used the mean square error (MSE) (Equation 11) as the cost function:
@@
Here, $Y_{i}$ and $\hat{Y}_{i}$ denote the predicted and target output variables, where $N$ is the number of observations. There was no need to manually set a threshold for the cost function, as BR automatically adjusts parameters like regularization terms and stopping criteria based on the data and model complexity.
@@
Figure 5: Generalizability of ANN models to new cylinder-based and sinusoid-based designs: (a-d) SEM images of cylinder-based fractal designs created by incorporating concentric rings, with corresponding unit cells shown below: (a) $C_{1}$, (b) $C_{2}$, (c) $C_{3}$, and (d) $C_{4}$. (e-i) SEM images of sinusoid-based designs with corresponding sine function, thickness and diameter of the inscribed cylinder given below: (e) $S_{1}$, (f) $S_{2}$, (g) $S_{3}$, (h) $S_{4}$, and (i) $S_{5}$. Representative specific stress-strain profiles, with the shaded regions indicating the standard deviation in the measured specific stress for (j) cylinder-based designs, (k) one-way sinusoid based designs and (l) two-way sinusoid-based designs. Comparison of values predicted by ANN with the experimentally measured values across architectures, $C_{1}$ to $S_{5}$ for (m) $W^{*}$, (n) $\sigma^{*}$ and (o) $E^{*}$. (p) Ratio ($\eta$) of ANN-predicted to the experimentally-measured values for $W^{*}$, $\sigma^{*}$ and $E^{*}$ across $C_{1}$ to $S_{5}$
@@
For each mechanical property, the ANN differed only in the optimal number of neurons in the hidden layer. The grid search-based parameter optimization process (Figure S1) that we employed is illustrated in Supplementary Information . We split the input dataset of 81 designs into two subsets, 66 designs for training, and 15 designs for testing. Using a holdout validation approach (shuffle-split procedure) [46, 44], we trained the network multiple times with different shuffled splits, then averaged the performance metrics to obtain a more reliable estimate of the model’s overall performance.
The ANN model demonstrated excellent correlations for all three mechanical properties under consideration, indicating its effectiveness in capturing the structure-property relationships. As shown in Fig. 4(b-g), the coefficient of determination ($R^{2}$) greater than 0.8 and mean absolute percentage errors (MAPE) around 10-15 % (Table 2) for each property highlight the model’s impressive performance after training. These results underscore the capability of the implicit geometric descriptor-based ANN approach to successfully establish a unified structure-property relationship that integrates various types of 2D architectures. We compared the performance of our ANN models with other regression models (see Supplementary Information), including multivariate linear regression (MVLR), MVLR with an added interdependence term given by the product of the descriptors (MVLR-I) (Equation S1) and Gaussian process regression (GPR) (Equation S2). Table 2 provides the summary of the $R^{2}$ and MAPE obtained with each model and demonstrates that the ANN models outperform other regression models in establishing accurate correlations. The relatively strong performance of other regression models also highlight the effectiveness of the implicit geometric descriptor-based approach in capturing the hidden structure-property correlations. Our ANN’s ability to establish a unified structure-property relationship using implicit geometric descriptors that capture the size effect and interaction effect suggests that, in architected VACNT foams, mechanisms like size effect and interaction effect are the primary drivers of bulk mechanical properties, irrespective of the specific 2D architecture—a finding that underscores the utility of the geometric descriptors in ANN rather than mere image regression-based approaches.
@@
Figure 6: Design of architected materials with implicit geometric descriptors. (a) Schematic illustration of the various training conditions used to evaluate the descriptor importance scores—$\Omega_{\mathcal{I}_{mc-c}}$ and $\Omega_{\mathcal{D}}$—and the coefficients of determination—$R^{2}_{\mathcal{I}_{mc-c}}$ and $R^{2}_{\mathcal{D}}$—for the descriptors $\mathcal{I}_{mc-c}$ and $\mathcal{D}$, respectively. (b) Comparison of the descriptor importance scores—$\Omega_{\mathcal{I}_{mc-c}}$ and $\Omega_{\mathcal{D}}$—demonstrating the influence of each descriptor on specific mechanical properties $W^{*}$, $\sigma^{*}$, and $E^{*}$; lower values of $\Omega_{\mathcal{D}}$ in spite of $\Omega_{\mathcal{I}_{mc-c}}$ not changing significantly implies the interdependence between the descriptors in influencing $\sigma^{*}$ and $E^{*}$. (c) Comparison of the coefficients of determination—$R^{2}_{\mathcal{I}_{mc-c}}$ and $R^{2}_{\mathcal{D}}$—showing similar trend as in the case of descriptor importance scores. Design contours for the following specific mechanical properties as functions of implicit geometric descriptors predicted by the ANN model: (d) $W^{*}$, (e) $\sigma^{*}$ and (f) $E^{*}$